what metal is the part with a mass of 7.8 kg and a volume of 1000 cm3 made of


1
Expert's answer
2021-09-21T06:33:54-0400

Given:

"V=1000\:\rm cm^3=1*10^{-3}\: m^3"

"m=7.8\:\rm kg"


The mass density

"\rho=\frac{m}{V}=\rm\frac{7.8\: kg}{1*10^{-3}\: m^3}=7800\: kg/m^3"

Answer: iron